# macOSWorld: A Multilingual Interactive Benchmark for GUI Agents

### v3

@@ Abstract
Graphical User Interface (GUI) agents show promising capabilities for automating computer-use tasks and facilitating accessibility, but existing interactive benchmarks are mostly English-only, covering web-use or Windows, Linux, and Android environments, but not macOS. macOS is a major OS with distinctive GUI patterns and exclusive applications. To bridge the gaps, we present macOSWorld, the first comprehensive benchmark for evaluating GUI agents on macOS. macOSWorld features 202 multilingual interactive tasks across 30 applications (28 macOS-exclusive), with task instructions and OS interfaces offered in 5 languages (English, Chinese, Arabic, Japanese, and Russian). As GUI agents are shown to be vulnerable to deception attacks, macOSWorld also includes a dedicated safety benchmarking subset. Our evaluation on six GUI agents reveals a dramatic gap: proprietary computer-use agents lead at above 30% success rate, while open-source lightweight research models lag at below 5%, highlighting the need for macOS domain adaptation. Multilingual benchmarks also expose common weaknesses, especially in Arabic, with a 28.8% average degradation compared to English. Results from safety benchmarking also highlight that deception attacks are more general and demand immediate attention. macOSWorld is available at https://github.com/showlab/macosworld.
@@ 1 Introduction
The benchmarking results reveal distinct performance tiers, with proprietary computer use agents (CUAs) achieving over 30% success rate while open-source research models struggle at below 5%. On average, the agents perform the best with linear alphabetic languages (English, Russian), followed by block character-based languages (Chinese, Japanese), while right-to-left Arabic shows a 28.8% performance drop compared to English. This drop primarily manifests through degraded planning or grounding capabilities. Task and environment language mismatch further degrades agent performance. Our safety evaluation reveals that agents’ susceptibility varies, with the two strongest CUAs being highly vulnerable to context deception attacks ($\approx$70deception rate), highlighting that the issue is more general than previously understood and demands immediate attention.
@@ 2 Related Works
OSWorld [19] 369+43 Ubuntu, Windows 1$\times$1 1 / 9
WindowsAgentArena [20] 154 Windows 1$\times$1 7 / 12
AndroidWorld [21] 116 Android 1$\times$1 18 / 20
WorldGUI [22] 107 Windows 1$\times$1 2 / 9
macOSWorld (Ours) 201+29 macOS 5$\times$5 28 / 30
@@
Despite significant advances, existing benchmarks omit macOS and its unique applications, lack coverage of non-English tasks or interfaces, and rarely combine functional performance with safety evaluation. We therefore develop macOSWorld, a multilingual interactive macOS benchmark integrated with safety evaluations to bridge the gaps.
@@
Figure 2: macOSWorld benchmark infrastructure. The main components are (1) a suite of multilingual tasks with natural-language instructions and programmatic evaluation, (2) interactive, reproducible macOS computer environments hosted on AWS, and (3) a centralized testbench that drives the evaluation process, orchestrates different components via SSH, VNC, and AWS APIs.
@@ 3.4 Testbench
macOSWorld comprises 202 interactive tasks organized into seven categories, of which 171 are available in five languages. These tasks span common system interfaces (such as the Lock Screen and App Launcher) as well as 30 macOS applications – 28 of which are exclusive to macOS. By benchmarking both macOS-unique applications and multilingual task execution (environment language and instruction language), macOSWorld fills two important gaps in prior work.
@@ 4.1 Task Instruction Curation
The task instructions were authored in English by our annotators according to the following principles. First, we adopted the task taxonomy from OSWorld [19] and WindowsAgentArena [20], which has been established through extensive computer use cases (see Table 9 in [19]). Second, we replaced Windows and Linux contexts with their macOS counterparts – for example, swapping LibreOffice for iWork, VSCode for Xcode, and Python coding tasks for SwiftUI development. Third, we reduced the proportion of web browsing scenarios (already well covered by benchmarks such as [13, 14, 19]) and instead emphasized interactions with macOS-unique application interfaces. Finally, annotators consulted official Apple resources – including tutorials, sample projects and templates – to ensure that each instruction reflected representative use cases and user flows. The instructions were then translated by GPT-4o into Chinese, Arabic, Japanese, and Russian versions (retaining proper nouns and file names), with Google Translate round-trip verification back to English to confirm consistency. Figure 5 provides an example of a task in all 5 languages, and screenshots of their initial states.
@@ 4.3 Safety Benchmarking Subset
Figure 4: Examples of our safety benchmarking subset. (a) Pop-up window attack [42] (top) versus our macOS-style deceptive pop-up window (bottom). (b) Four examples of our attack. Our method spawns real pop-up windows in the environment with several buttons. Only when the distracting buttons are clicked would the attack be considered successful.
@@
We build on the pop-up window attack of prior work [42], adapting it to macOS and making three key modifications: (1) UI Consistency: We replaced the original’s exaggerated typography with fonts and sizes faithful to standard macOS dialogs. (2) Evaluation Criteria: Each popup has both "gold" and "distraction" buttons, and only clicking the distraction buttons counts as a successful attack. (3) Element Obstruction: Dialogs are centered on the screen, potentially obscuring critical UI elements and requiring explicit handling. Figure 4(a) compares [42]’s synthetic attack (top) with our version (bottom).
@@ 5 Benchmarking Baselines
Model & Language Success Rate ($\uparrow$)
@@ 5.1 Benchmark Setup
We evaluate six representative GUI agents as baselines: two proprietary computer-use agents (OpenAI Computer-Using Agent [28], computer-use-preview-2025-03-11; Claude Computer-Use Agent [29], claude-3-7-sonnet-20250219 with computer-use-2025-01-24 and token-efficient-tools-2025-02-19 betas), two general VLM-based agents (GPT-4o [46], gpt-4o-2024-08-06; Gemini 2.5 Pro [47], gemini-2.5-pro-preview-03-25), and two community open-source GUI agents (ShowUI 2B [25]; UI-TARS 7B DPO [2], chain-of-thought [48] enabled in Chinese). This selection spans state-of-the-art proprietary systems, powerful vision-language backbones, and lightweight research models. Implementation details (including prompts used, temperature, and top_p) are given in Appendix G.
@@
All experiments run in a macOS virtual instance at 1024$\times$768 pixels (the default for macOS and for Claude CUA [29]), physically on AWS-hosted Mac Minis (model A2348). Except for the Set-of-Mark (SoM) ablation (Table 8), all agents perceive the environment from the most recent 3 screenshot only. Each task is granted up to 15 screenshots or 30 dialog turns, whichever limit is reached first. Agents retain the full conversation history, but prune screenshots to the most recent 3 (with ShowUI [25] being an exception, following its own context format [4]). Tasks are benchmarked in five languages – English (en), Chinese (zh), Arabic (ar), Japanese (jp), and Russian (ru) – with both the task prompt and system UI set to the target language. The only exception is the Advanced Apps category, which is evaluated in English exclusively (Section 4.2).
@@ 5.2 Quantitative Performance Evaluation
As shown in Table 3), the two proprietary CUAs lead with overall SRs above 30%. Gemini 2.5 Pro, alone, occupy the middle tier, at 18.5%. Finally, GPT-4o and the two open-source models (UI-TARS 7B DPO, ShowUI) register below 10% on average. For the open-source models, compared to their competitive performance on web browsing, Windows and Ubuntu Linux [25, 2, 19], the results here reveal they lack macOS-specific adaptation. Notably, Although GPT-4o and UI-TARS are close in performance, their failure modes are completely different, which we analyze in detail in Section 5.3.
Agents handle interface and productivity tasks moderately but struggle on media, advanced apps, and multi-app workflows.
When averaged over the four interface and productivity categories (System & Interface, System Apps, File Management, Productivity), agents achieve a fairly consistent 17%-21% SR, demonstrating basic navigation and command execution competence. However, Media tasks fall to 12.8% SR – likely because these involve more precise dragging operations that current agents handle poorly. Advanced Apps tasks, which demand both domain knowledge and complex operations, peak at only 19.4% (OpenAI CUA) and drop to 0% for even GPT-4o. Multi-app scenarios are the hardest of all, with a mere 3.7% average SR, underscoring the challenge of coordinating actions across concurrent windows and applications.
@@
As shown in Table 4, across five languages evaluated, linear alphabetic languages (English, Russian) yield the highest average overall SRs across agents (>17.5%). East Asian block characters (Chinese, Japanese) follow closely at 17.2% and 15.8%. Right-to-left Arabic lags at 13.7%, which is a 28.8% drop from English, reflecting agents’ difficulty perceiving either the visual complexity of Arabic glyphs or the mirrored UI layout, or both. GPT-4o exhibits especially uneven multilingual performance, with over 60% SR degradation on Russian and Arabic compared to English. If the task language and the environment language do not match, agent performance could further decline (see Appendix D.1).
@@ 5.3 Qualitative Analysis: Agent Behavior and Failure Modes
ShowUI demonstrates decent grounding ability but often executes nonsensical actions. For example, it attempts to open Settings by scrolling the mouse wheel on the desktop (Figure 9). It also tries to complete the task by typing it into the Help menu (Figure 8), which clearly shows that while it can identify basic interface elements, it lacks macOS-specific domain knowledge. In comparison, UI-TARS not only exhibits similarly nonsensical behaviors (Figure 10 and 14), but also fails due to illegal output formats and hallucinations. In Figure 15, it issues instructions to click a button by name rather than by coordinates, which is an illegal operation; in Figure 14, it mistakenly believes the Apple menu is located at the bottom-left corner of the screen. These results indicate that lightweight research models require substantial adaptation for mainstream macOS GUI tasks.
@@
Although GPT-4o’s performance is close to UI-TARS, its failure mode is more similar to Gemini 2.5 Pro, another general-purpose VLM. Both GPT and Gemini struggle to precisely ground GUI elements, consistent with prior findings on general VLM agents in GUI settings [19, 20, 31]. While Gemini 2.5 Pro shows noticeably better grounding than GPT-4o, neither approaches the performance of specialized computer-use agents. Grounding UI elements with Set-of-Mark annotations [31] doubles GPT-4o’s performance (details in Appendix D.2), at 13.0% averaged across all languages (Table 8), but still lags Gemini’s 18.5%.
@@
Differences in performance across languages are most pronounced for the two proprietary CUAs. For OpenAI CUA, the performance degradation stem primarily from degraded grounding: In one task (Figure 18), the agent effortlessly completed the account-creation task in English, but in Arabic, it fails to open Settings at all. Claude CUA exhibits both grounding and planning-related errors: while GUIs are mirrored in Arabic, Claude may carry biases from left-to-right UI layouts and incorrectly click on non-mirrored locations. It could also use more steps to complete the same task (Figure 19). These cross-language discrepancies underscore the necessity for language-aware training and UI layout adaptation to maintain consistent performance across diverse language environments.
@@ 5.4 Safety Subset Evaluation
Table 5: Agent performance on the macOSWorld safety subset under context-deception attacks, showing rates (%) of Distracted (clicking a decoy), Gold (clicking the close/cancel button), and Unhandled (pop-up not closed). Colors indicate highest and second highest rates in each row.
@@ Appendix B Ethical and Safeguarding Statement
For responsible deployment, macOSWorld implements environment isolation for safeguarding. The benchmark operates solely within virtualized macOS environments on AWS Mac mini machines, isolating potential harmful operations from both host platforms and physical hardware. Our snapshot restoration process ensures all content, files, and system changes are immediately discarded after each evaluation, preventing any accumulation of sensitive information.
@@ Appendix C Example of a Real Task
Before the agent begins, we must restore the macOS environment to the exact state in which the target document is already open. We do this by launching a pre-configured Amazon Machine Image (AMI) on AWS EC2. Each language variant maps to its own AMI ID. For example: { "en": "ami-07f4fd69378358c18", "zh": "ami-0abc1234def567890", ... }.
@@
Once the EC2 instance is up, our testbench runs a language-specific preparation script over SSH (via zsh and AppleScript). The script given in Figure 5 (1) launches Numbers, (2) waits five seconds to avoid race conditions, and (3) opens a new document from the built-in "Personal Budget" template. Because template names differ across macOS languages, for some tasks, we maintain one script per language.
@@
Although this example uses a single evaluation script, many macOSWorld tasks employ multiple scripts to handle language-dependent window titles or alternative success criteria. While our platform supports non-binary rewards for intermediate states, the experiments reported in this paper remove those rewards and use only binary (success/failure) evaluations.
@@ D.1 Cross-Lingual Evaluation
Language System & Interface System Apps File Ops Produc- tivity Media Multi- tasking Total
@@
Task Language Environment Language
@@ D.2 Set-of-Mark (SoM) Annotation Evaluation
Table 8 compares GPT-4o’s SR across the five languages when using explicit SoM tag identifiers in place of raw screen coordinates. The results show that SoM annotations roughly double GPT-4o’s success rates in every language. Nevertheless, GPT-4o with SoM still trails behind Gemini Pro 2.5 and both proprietary CUAs. This gap indicates that SoM annotations are a useful but insufficient workaround; agents specifically trained or finetuned for computer-use tasks maintain superior grounding and overall performance.

### v4

@@ Abstract
Graphical User Interface (GUI) agents show promising capabilities for automating computer-use tasks and facilitating accessibility, but existing interactive benchmarks are mostly English-only, covering web-use or Windows, Linux, and Android environments, but not macOS. macOS is a major OS with distinctive GUI patterns and exclusive applications. To bridge the gaps, we present macOSWorld, the first comprehensive benchmark for evaluating GUI agents on macOS. macOSWorld features 202 multilingual interactive tasks across 30 applications (28 macOS-exclusive), with task instructions and OS interfaces offered in 5 languages (English, Chinese, Arabic, Japanese, and Russian). As GUI agents are shown to be vulnerable to deception attacks, macOSWorld also includes a dedicated safety benchmarking subset. Our evaluation on six GUI agents reveals a dramatic gap: proprietary computer-use agents lead at above 30% success rate, while open-source lightweight research models lag at below 5%, highlighting the need for macOS domain adaptation. Multilingual benchmarks also expose common weaknesses, especially in Arabic, with a 28.8% average degradation compared to English. Results from safety benchmarking also highlight that deception attacks are more general and demand immediate attention. Project page: https://macos-world.github.io.
@@ 1 Introduction
The benchmarking results reveal distinct performance tiers, with proprietary computer use agents (CUAs) achieving over 30% success rate while open-source research models struggle at below 5%. On average, the agents perform the best with linear alphabetic languages (English, Russian), followed by block character-based languages (Chinese, Japanese), while right-to-left Arabic shows a 28.8% performance drop compared to English. This drop primarily manifests through degraded planning or grounding capabilities. Task and environment language mismatch further degrades agent performance. Our safety evaluation reveals that agents’ susceptibility varies, with the two strongest CUAs being highly vulnerable to context deception attacks ($\approx$70% deception rate), highlighting that the issue is more general than previously understood and demands immediate attention.
@@ 2 Related Works
OSWorld [19] 369+43 Ubuntu, Windows $\times1$ 1 / 9
WindowsAgentArena [20] 154 Windows $\times1$ 7 / 12
AndroidWorld [21] 116 Android $\times1$ 18 / 20
WorldGUI [22] 107 Windows $\times1$ 2 / 9
macOSWorld (Ours) 201+29 macOS $\times5$ 28 / 30
@@
Despite significant advances, existing benchmarks omit macOS and its unique applications, lack coverage of non‑English tasks or interfaces, and rarely combine functional performance with safety evaluation. We therefore develop macOSWorld, a multilingual interactive macOS benchmark integrated with safety evaluations to bridge the gaps.
@@
Figure 2: macOSWorld benchmark infrastructure. The main components are (1) a suite of multilingual tasks with natural‐language instructions and programmatic evaluation, (2) interactive, reproducible macOS computer environments hosted on AWS, and (3) a centralized testbench that drives the evaluation process, orchestrates different components via SSH, VNC, and AWS APIs.
@@ 3.4 Testbench
macOSWorld comprises 202 interactive tasks organized into seven categories, of which 171 are available in five languages. These tasks span common system interfaces (such as the Lock Screen and App Launcher) as well as 30 macOS applications – 28 of which are exclusive to macOS. By benchmarking both macOS‑unique applications and multilingual task execution (environment language and instruction language), macOSWorld fills two important gaps in prior work.
@@ 4.1 Task Instruction Curation
The task instructions were authored in English by our annotators according to the following principles. First, we adopted the task taxonomy from OSWorld [19] and WindowsAgentArena [20], which has been established through extensive computer use cases (see Table 9 in [19]). Second, we replaced Windows and Linux contexts with their macOS counterparts – for example, swapping LibreOffice for iWork, VSCode for Xcode, and Python coding tasks for SwiftUI development. Third, we reduced the proportion of web browsing scenarios (already well covered by benchmarks such as [13, 14, 19]) and instead emphasized interactions with macOS‑unique application interfaces. Finally, annotators consulted official Apple resources – including tutorials, sample projects and templates – to ensure that each instruction reflected representative use cases and user flows. The instructions were then translated by GPT‑4o into Chinese, Arabic, Japanese, and Russian versions (retaining proper nouns and file names), with Google Translate round‑trip verification back to English to confirm consistency. Figure 5 provides an example of a task in all 5 languages, and screenshots of their initial states.
@@ 4.3 Safety Benchmarking Subset
Figure 4: Examples of our safety benchmarking subset. (a) Pop-up window attack [42] (top) versus our macOS‑style deceptive pop-up window (bottom). (b) Four examples of our attack. Our method spawns real pop-up windows in the environment with several buttons. Only when the distracting buttons are clicked would the attack be considered successful.
@@
We build on the pop‑up window attack of prior work [42], adapting it to macOS and making three key modifications: (1) UI Consistency: We replaced the original’s exaggerated typography with fonts and sizes faithful to standard macOS dialogs. (2) Evaluation Criteria: Each popup has both "gold" and "distraction" buttons, and only clicking the distraction buttons counts as a successful attack. (3) Element Obstruction: Dialogs are centered on the screen, potentially obscuring critical UI elements and requiring explicit handling. Figure 4(a) compares [42]’s synthetic attack (top) with our version (bottom).
@@ 5 Benchmarking Baselines
Model &
Language
Success Rate ($\uparrow$)
@@ 5.1 Benchmark Setup
We evaluate six representative GUI agents as baselines: two proprietary computer‑use agents (OpenAI Computer‑Using Agent [28], computer-use-preview-2025-03-11; Claude Computer‑Use Agent [29], claude-3-7-sonnet-20250219 with computer-use-2025-01-24 and token-efficient-tools-2025-02-19 betas), two general VLM‑based agents (GPT‑4o [46], gpt-4o-2024-08-06; Gemini 2.5 Pro [47], gemini-2.5-pro-preview-03-25), and two community open‑source GUI agents (ShowUI 2B [25]; UI‑TARS 7B DPO [2], chain-of-thought [48] enabled in Chinese). This selection spans state‑of‑the‑art proprietary systems, powerful vision‑language backbones, and lightweight research models. Implementation details (including prompts used, temperature, and top_p) are given in Appendix G.
@@
[…]
@@ 5.2 Quantitative Performance Evaluation
As shown in Table 3), the two proprietary CUAs lead with overall SRs above 30%. Gemini 2.5 Pro, alone, occupy the middle tier, at 18.5%. Finally, GPT-4o and the two open-source models (UI‑TARS 7B DPO, ShowUI) register below 10% on average. For the open-source models, compared to their competitive performance on web browsing, Windows and Ubuntu Linux [25, 2, 19], the results here reveal they lack macOS-specific adaptation. Notably, Although GPT-4o and UI-TARS are close in performance, their failure modes are completely different, which we analyze in detail in Section 5.3.
Agents handle interface and productivity tasks moderately but struggle on media, advanced apps, and multi‑app workflows.
When averaged over the four interface and productivity categories (System & Interface, System Apps, File Management, Productivity), agents achieve a fairly consistent 17%-21% SR, demonstrating basic navigation and command execution competence. However, Media tasks fall to 12.8% SR – likely because these involve more precise dragging operations that current agents handle poorly. Advanced Apps tasks, which demand both domain knowledge and complex operations, peak at only 19.4% (OpenAI CUA) and drop to 0% for even GPT-4o. Multi‑app scenarios are the hardest of all, with a mere 3.7% average SR, underscoring the challenge of coordinating actions across concurrent windows and applications.
@@
As shown in Table 4, across five languages evaluated, linear alphabetic languages (English, Russian) yield the highest average overall SRs across agents (>17.5%). East Asian block characters (Chinese, Japanese) follow closely at 17.2% and 15.8%. Right‑to‑left Arabic lags at 13.7%, which is a 28.8% drop from English, reflecting agents’ difficulty perceiving either the visual complexity of Arabic glyphs or the mirrored UI layout, or both. GPT‑4o exhibits especially uneven multilingual performance, with over 60% SR degradation on Russian and Arabic compared to English. If the task language and the environment language do not match, agent performance could further decline (see Appendix D.1).
@@ 5.3 Qualitative Analysis: Agent Behavior and Failure Modes
ShowUI demonstrates decent grounding ability but often executes nonsensical actions. For example, it attempts to open Settings by scrolling the mouse wheel on the desktop (Figure 9). It also tries to complete the task by typing it into the Help menu (Figure 8), which clearly shows that while it can identify basic interface elements, it lacks macOS‑specific domain knowledge. In comparison, UI‑TARS not only exhibits similarly nonsensical behaviors (Figure 10 and 14), but also fails due to illegal output formats and hallucinations. In Figure 15, it issues instructions to click a button by name rather than by coordinates, which is an illegal operation; in Figure 14, it mistakenly believes the Apple menu is located at the bottom-left corner of the screen. These results indicate that lightweight research models require substantial adaptation for mainstream macOS GUI tasks.
@@
Although GPT-4o’s performance is close to UI-TARS, its failure mode is more similar to Gemini 2.5 Pro, another general-purpose VLM. Both GPT and Gemini struggle to precisely ground GUI elements, consistent with prior findings on general VLM agents in GUI settings [19, 20, 31]. While Gemini 2.5 Pro shows noticeably better grounding than GPT‑4o, neither approaches the performance of specialized computer‑use agents. Grounding UI elements with Set-of-Mark annotations [31] doubles GPT-4o’s performance (details in Appendix D.2), at 13.0% averaged across all languages (Table 8), but still lags Gemini’s 18.5%.
@@
Differences in performance across languages are most pronounced for the two proprietary CUAs. For OpenAI CUA, the performance degradation stem primarily from degraded grounding: In one task (Figure 18), the agent effortlessly completed the account-creation task in English, but in Arabic, it fails to open Settings at all. Claude CUA exhibits both grounding and planning‑related errors: while GUIs are mirrored in Arabic, Claude may carry biases from left-to-right UI layouts and incorrectly click on non-mirrored locations. It could also use more steps to complete the same task (Figure 19). These cross‑language discrepancies underscore the necessity for language‑aware training and UI layout adaptation to maintain consistent performance across diverse language environments.
@@ 5.4 Safety Subset Evaluation
Table 5: Agent performance on the macOSWorld safety subset under context‑deception attacks, showing rates (%) of Distracted (clicking a decoy), Gold (clicking the close/cancel button), and Unhandled (pop-up not closed). Colors indicate highest and second highest rates in each row.
@@ 7 Acknowledgements
This research is supported by the National Research Foundation, Singapore under its AI Singapore Programme (AISG Award No: AISG3-RP-2022-030).
@@ Appendix B Ethical and Safeguarding Statement
For responsible deployment, macOSWorld implements environment isolation for safeguarding. The benchmark operates solely within virtualized macOS environments on AWS Mac mini machines, isolating potential harmful operations from both host platforms and physical hardware. Our snapshot restoration process ensures all content, files, and system changes are immediately discarded after each evaluation, preventing any accumulation of sensitive information. To further facilitate auditability and traceability, additional techniques such as output watermarking could be employed [49, 50, 51, 52].
@@ Appendix C Example of a Real Task
Before the agent begins, we must restore the macOS environment to the exact state in which the target document is already open. We do this by launching a pre‑configured Amazon Machine Image (AMI) on AWS EC2. Each language variant maps to its own AMI ID. For example: { "en": "ami-07f4fd69378358c18", "zh": "ami-0abc1234def567890", ... }.
@@
Once the EC2 instance is up, our testbench runs a language‑specific preparation script over SSH (via zsh and AppleScript). The script given in Figure 5 (1) launches Numbers, (2) waits five seconds to avoid race conditions, and (3) opens a new document from the built‑in "Personal Budget" template. Because template names differ across macOS languages, for some tasks, we maintain one script per language.
@@
Although this example uses a single evaluation script, many macOSWorld tasks employ multiple scripts to handle language‑dependent window titles or alternative success criteria. While our platform supports non‑binary rewards for intermediate states, the experiments reported in this paper remove those rewards and use only binary (success/failure) evaluations.
@@ D.1 Cross-Lingual Evaluation
Language
System &
Interface
System
Apps
File
Ops
Produc-
tivity
Media
Multi-
tasking
Total
@@
Task
Language
Environment Language
@@ D.2 Set-of-Mark (SoM) Annotation Evaluation
Table 8 compares GPT-4o’s SR across the five languages when using explicit SoM tag identifiers in place of raw screen coordinates. The results show that SoM annotations roughly double GPT-4o’s success rates in every language. Nevertheless, GPT-4o with SoM still trails behind Gemini Pro 2.5 and both proprietary CUAs. This gap indicates that SoM annotations are a useful but insufficient workaround; agents specifically trained or finetuned for computer‑use tasks maintain superior grounding and overall performance.